40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 25 15:43:29 2016, ericq, Update, SUS, OSEM issues SRM_OSEMS.pngPRM_LR.png
    Reply  Thu Aug 25 20:07:35 2016, gautam, Update, SUS, OSEM issues - maybe resolved? 
    Reply  Tue Sep 6 18:21:13 2016, rana, Update, SUS, OSEM issues PRMswing.png
       Reply  Tue Sep 6 20:52:42 2016, gautam, Update, SUS, OSEM issues P9060254.JPG
          Reply  Wed Sep 7 17:23:26 2016, gautam, Update, SUS, PRM LR fixed for now IMG_3208.JPGIMG_3211.JPG
             Reply  Tue Oct 11 00:43:50 2016, ericq, Update, SUS, PRM LR problematic again PRM_LR_Problem.pngsince_fix.pnglast10.png
                Reply  Tue Oct 11 10:15:04 2016, Steve, Update, SUS, PRM LR problematic again PRMalfoiled.png
                   Reply  Tue Oct 11 13:30:49 2016, gautam, Update, SUS, PRM LR problematic again PRM_LR_decay.png
                      Reply  Wed Oct 12 13:34:28 2016, gautam, Update, SUS, PRM LR problematic again 08.png
Message ID: 12443     Entry time: Thu Aug 25 20:07:35 2016     In reply to: 12441
Author: gautam 
Type: Update 
Category: SUS 
Subject: OSEM issues - maybe resolved? 

[lydia, johannes, gautam]

While struggling to minimize the bounce mode coupling into the sensor signals, we briefly poked into the ITMY chamber, and think that we understand the origin of the problem, at least for the SRM.

Essentially, we believe that moving the ITM from its nominal position to the edge of the table has shifted the table leveling such that the optic (SRM) is tilted backwards (hence the magnets are completely occluding the LEDs) and that perhaps the optic is in contact with one or more of the bottom EQ stops (hence the signal is stationary, no oscillations visible. The timing of the signals going dark as Eric mentioned supports this hypothesis. The reason why we believe this to be the case is that when I was trying to loosen the screw on the clamp holding the ITMY cage to the table, we saw ~1Hz signals from all 5 SRM OSEM sensors, though they were well away from the nominal equilibrium values. The arrangement of towers in the chamber right now did not permit me to get a good look at the SRM magnets, but I believe they are all still attached to the optic, and that they are NOT stuck to the OSEM coils. If this is indeed the case, putting ITMY back in will solve the issue completely.

It is not clear what has happened to the LR coil on the PRM - could it be that during the venting process, somehow the LR magnet got stuck to the OSEM? If so, can we free it by the usual bias jiggling?

ELOG V3.1.3-